The town recreation department ordered a total of 100 balls and bats and hats for the summer baseball camp. Balls cost $4.50 each, and bats cost $20.00 each. The total purchase was $822. How many of each item were ordered?
Answer by stanbon(75887)

The town recreation department ordered a total of 100 balls and bats and hats for the summer baseball camp. Balls cost $4.50 each, and bats cost $20.00 each. The total purchase was $822. How many of each item were ordered?
--------
Equations:
b + h = 100
4.5b + 20h = 822
------
Multiply thru 1st by 200
Multiply thru 2nd by 10.
--------
200b + 200h = 200*100
45b + 200h = 8220
------------------------
Subtract and solve for "b":
155b = 11780
# of balls = 76
----
# of bats = 100-76 = 24
--------------
Cheers,
Stan H.
===================